# GSQA: An End-to-End Model for Generative Spoken Question Answering

### 15 Dec 2023 v1

@@ Abstract
In recent advancements in spoken question answering (QA), end-to-end models have made significant strides. However, previous research has primarily focused on extractive span selection. While this extractive-based approach is effective when answers are present directly within the input, it falls short in addressing abstractive questions, where answers are not directly extracted but inferred from the given information. To bridge this gap, we introduce the first end-to-end Generative Spoken Question Answering (GSQA) model that empowers the system to engage in abstractive reasoning. The challenge in training our GSQA model lies in the absence of a spoken abstractive QA dataset. We propose using text models for initialization and leveraging the extractive QA dataset to transfer knowledge from the text generative model to the spoken generative model. Experimental results indicate that our model surpasses the previous extractive model by 3% on extractive QA datasets. Furthermore, the GSQA model has only been fine-tuned on the spoken extractive QA dataset. Despite not having seen any spoken abstractive QA data, it can still closely match the performance of the cascade model. In conclusion, our GSQA model shows the potential to generalize to a broad spectrum of questions, thus further expanding spoken question answering capabilities of abstractive QA. Our code is available at https://voidful.github.io/GSQA
Index Terms—  Spoken Question Answering, Textless NLP, Generative Language models, Transfer learning
@@ 1 Introduction
An intuitive approach to tackle this challenge is the cascade method. Here, an Automatic Speech Recognition (ASR) system first transcribes the input speech of both questions and the passages, and then feeds transcriptions into a textual QA model that predicts the corresponding answers. Last, answers are transformed back into speech through a Text-To-Speech (TTS) model. The cascade method’s primary advantage is its ability to harness the power of text-based QA models trained on vast datasets, ensuring it can tackle both extractive and abstractive QA tasks. However, the cascade model has its drawbacks. The primary issue stems from error propagation, since the text-based QA model isn’t typically trained on ASR-erroneous data, the presence of ASR errors can notably undermine the performance of cascade models.
@@
Recently, end-to-end, textless approaches (e.g., DUAL[5]) have emerged as a solution to the cascade method’s limitations. These works [5, 6, 7] rely on the representations of acoustic models, like HuBERT[8], and apply K-means clustering to transform speech into discrete units. These units, encapsulating both questions and passages, are fed into the model that predicts the span of the answer in the passage using start and end positions. By adopting this textless approach, DUAL effectively sidesteps the error propagation from ASR systems [9, 10, 11]. It improves the model robustness in noisy speech. However, there’s a notable limitation to the DUAL model: it is not able to handle abstractive QA scenarios as its design to only output start and end positions within the input.
As Fig.1 shows the different of our purpose method, cascade method, and DUAL, we combine the advantages of both the cascade model and the end-to-end textless approach to introduce the first-ever textless end-to-end model capable of handling both extractive and abstractive QA. Mirroring the DUAL’s methodology, we utilize HuBERT to convert the input speech (question and passage) into discrete units[6]. A sequence-to-sequence model then processes these units to generate the answer in discrete units. Motivated by the strengths and limitations of approaches above, we propose GSQA (Generative Spoken Question Answering), the first textless end-to-end model capable of handling both extractive and abstractive QA.
Furthermore, given the current absence of a spoken abstractive QA corpus, we begin by initializing GSQA with a generative Text QA (TQA) model. This TQA model has been trained on several extractive and abstractive datasets using a generative approach. Subsequently, we fine-tune the TQA model on NMSQA, a dataset dedicated to extractive spoken QA. GSQA primary objective is to impart the Natural Language Understanding prowess of the TQA model to GSQA. We also use Microsoft’s Azure TTS service to create a synthesized test set for evaluation. It’s noteworthy that, during the transfer learning phase, GSQA exhibited a significant enhancement, registering an 18% relative gain in the BLEU1 score in a zero-shot abstractive spoken QA setting.
@@
    Demonstrating that our generative QA model has competitive performance against textless extractive models.
@@ 2 Related Work
Most existing SQA methods rely on Automatic Speech Recognition (ASR) transcripts and then use a large language model (LLM) to fulfill the target of question-answering tasks. However, cascade models are strongly influenced by the ASR errors when an ASR system [12, 13] is not so robust that would pass wrong documents to LLM, and that causes LLM to predict wrong answers. On the other hand, a robust ASR system, like Whisper-large [13], which has 1550M parameters, can achieve a low word error rate, and can pass accurate transcriptions of documents to LLM to execute QA tasks well. However, though cascade models can achieve quite surprising accuracy on QA tasks when leveraging such an awesome ASR system, the large number of models’ parameters would take longer inference time.
Extractive SQA is to find the answer from a spoken document given a question. DUAL [5] is designed to predict the start and end index of answer spans. However, the limitation of DUAL is that it can only predict the answer span of the input audio, which means if answers do not exist in given spoken documents, it becomes useless.
TWIST [14] is an approach for solving the limitation of SpeechLM [15], which is the lack of semantic understanding that might lead to off-topic or wrong responses. However, though TWIST finds a better initialization point for SpeechLM, the experiment results of TWIST still do not show it has major semantic knowledge transfer.
@@ 3.1 Overview & Formulation
In the realm of Question Answering (QA), we divide the problem into three components: the Question (Q), the Passage (P), and the Answer (A). These entities can manifest in two formats: textual data ($Q_{t}$, $P_{t}$, $A_{t}$) and speech discrete unit data ($Q_{u}$, $P_{u}$, $A_{u}$).
Our method lies in the ambition to create a textless generative end-to-end QA model. Initially, we follow [6, 7] to quantize speech into discrete units. We then utilize a discrete unit-based sequence generation model to craft the answer units, which would be subsequently transformed into speech via a Unit-based HiFi-GAN vocoder [16]. To boost the logical understanding capabilities of the discrete unit-based sequence generation model, we initialize this generative spoken QA model using weights from a textual model [14]. Fig.2 shows the entire model architecture and training pipeline.
Fig. 2: Left: The process of discrete unit quantization from synthesis data. Right: Model Training Procedure: A depiction of the transition from textual QA pre-training to spoken QA fine-tuning.
@@ 3.2 Speech Quantization
The HuBERT[8] model serves as a formidable speech representation model. It collects a general representation by undertaking mask segment prediction training on its intermediate representation using k-means clustering. During this, every 20ms frame is clustered into one of the K categories. We obtain the k-means centroid id as the discrete representation of the frame, effectively transforming speech signals into discrete units. Following the methods presented in Lee et al. [6], we employed the HuBERT-base model and a 100 cluster number setting to achieve this transformation. Furthermore, to accelerate the process, we adopt the run-length encoding strategy to collapse consecutive sequences of identical units into a singular unit, thereby shortening the sequence length.
3.3 Text QA Model Pretraining
Our generative spoken QA model’s training data is relatively scarce, which could hamper the model’s performance. To tackle this issue, we suggest leveraging a pretrained textual model to amplify its semantic understanding. Moreover, this textual model is initially trained as a generative QA model before being transferred to a spoken QA model. This approach ensures that even if our spoken QA data is limited, the model can learn to transfer from text to speech units, allowing the capabilities of textual QA to be transferred onto the unit.
When converting speech to units, and even after deduplicating tokens, the resulting speech sequence remains longer than the text. Many generative textual models[17] can not handle such long input sequences. Therefore, when choosing our initialization textual model, it had to fulfill two criteria: 1) Be a generative model, and 2) Handle sequences exceeding 4096 in length[18, 19, 20].
In the end, we choose LongT5[18]. LongT5 extends the original T5 encoder, integrating global-local attention sparsity patterns to manage extended inputs. It marries the efficient attention mechanism of ETC with the summarization-centric pre-training objective of PEGASUS, resulting in exceptional performance. Upon this foundation, we further pretrain LongT5 on textual QA datasets.
The primary ambition during the pre-training phase is to develop a model capable of accepting a question $Q_{t}$ and passage $P_{t}$ in textual form and predicting the corresponding answer $A_{t}$. This training procedure is devised to equip our model with the essential capabilities needed for the ensuing stages of our method.
@@ 3.4 GSQA: Generative Spoken QA Model
Fig.2 illustrates our GSQA training procedure. Initially, we train the LongT5 on Text QA to produce the TQA LongT5 model, which is capable of text-level generative QA. This model is then fine-tuned on the Extractive Spoken QA dataset, NMSQA. We converted the NMSQA into a generative QA format, and it will output the answer directly in discrete unit form instead of the answer’s start and end position from extractive QA. Prior to this, both the input and output audio go through quantization to be converted into discrete units. We also employ a vocoder, based on the HiFi-GAN vocoder from [21], to convert these discrete units back to speech. This vocoder includes a duration prediction module for deduplicated unit synthesis. Since TQA-LongT5 doesnt́ recognize discrete units, we introduce them as new tokens. Their embeddings are initialized by random sampling from other text token embeddings, ensuring proper tokenization and a seamless transition from text to speech.
NMSQA-dev NMSQA-test Spoken-NarrativeQA-test Model Num. Param. F1 Score EM Score F1 Score EM Score BLEU1 ROUGE-L Cascade model (Upper bound) 1025M 49.1 32.0 47.3 30.4 13.5 19.9 DUAL 452M 39.4 21.9 33.6 21.2 - - Unit-LongT5 312M 25.5 12.6 20.1 9.4 6.8 10.4 Unit-LongT5-TQA (Proposed) 312M 41.8 24.9 36.0 24.0 8.0 11.8
Table 1: The table presents the Speech to Speech Evaluation on our models and the number of parameters for each method. For extractive QA datasets, we utilized F1 and exact match (EM) as evaluation metrics. On the other hand, for generative QA datasets, we employed BLEU1[22] and Rouge-L[23] for assessment. The cascade model in this table includes an ASR system and a LM, which is Whisper-medium.en [13] and LongT5-TQA[18], respectively. DUAL[5] is combined a HuBERT-base[8], K-means model for 128 clusters, and a Longformer[24].
@@ 4.1 Dataset
We use NMSQA dataset [5] as extractive SQA dataset. The NMSQA is derived from SQuAD v1.1, consists of Wikipedia paragraphs with human-written questions. While its train and dev sets feature synthesized speech using Amazon Polly, the test set contains audio from 60 human speakers (30 males and 30 females). In addition, we propose a new spoken abstractive SQA evaluation set, named Spoken-NarrativeQA. This dataset is built upon NarrativeQA test set, a abstractive text QA dataset for reading comprehension from books and movie scripts with complex narrative understanding. Note that many answers in this dataset are not directly extractable from passages. We transform the text to speech using TTS system and specifically select test data where answers are absent in the passage. There are 1626 testing samples in the Spoken-NarrativeQA dataset. For TQA training data, We utilized four extractive QA datasets: SQuAD [25], MultiRC [26], NewsQA [27], and Drop [28], along with one abstractive dataset, NarrativeQA [29], to train LM, and use the trained LM as foundation model of GSQA.
@@
Table 2: The respective number of parameters in our experiment. Note that E2E method, like DUAL and our work, does not need ASR model but need Unit-Extraction model to generate units, and the number after the name of that is the number of clusters.
@@
4.2 Cascaded Pipeline
The conventional approach for SQA tasks typically follows a cascaded pipeline, as illustrated in Figure 1 middle. The passage and question are first transcribed using ASR, and then a text-based QA model generates the textual answer. To generate the final spoken answer, we utilize the Azure TTS service to generate the spoken answer. This methodology serves as our baseline and is referred to as the Cascade Model. More specifically, our ASR model is using Whisper-medium.en [13], which has the state-of-the-art performance with comparable parameter size to DUAL and ours model. For the QA component, we employ the fine-tuned LongT5-TQA model, as mentioned in Section 3.3. The ASR performance of datasets is in Table 3.
4.3 Performance Evaluation
Since all our models ultimately output as a waveform, what we focus on in QA is whether the content of this waveform contains the answer. We then use an ASR model to convert the generated waveform back to text, allowing us to use text-based metrics to evaluate the content of the model. To reduce errors from the ASR during evaluation, we utilize the state-of-the-art ASR model, whisper-large-v2, during performance evaluation. We then compare the transcriptions of the outputs from whisper-large-v2 with the text ground truth. We calculate the F1-score and Exact-matched (EM) scores for extractive QA datasets. The BLEU score [22] and the Rouge score [23] are used for generative QA datasets. Lastly, the experimental results are reported in Table 1.
4.4 Implementation details
As the previous section mentioned, we pretrain LongT5 on the 5 TQA datasets for 13 epochs, and the learning rate is set to 0.0005, and weight decay is 0.01. On the other hand, when we fine-tune the pre-trained LongT5 on a unit-based dataset, NMSQA, for 25 epochs, the learning rate is set to 0.0003, and weight decay is 0.001. Last, when we inference our model on abstractive test set(e.g., Spoken-NarrativeQA-test), to encourage GSQA output longer length, we use beam search algorithm for decoding with a number of beam 5, and set length penalty to 2.
5 Results
As illustrated in Table 1, we evaluate on both extractive QA dataset NMSQA-dev, NMSQA-test, and abstractive QA dataset Spoken-Narrative QA. The following discussions provide insights into our main experiment results from different aspects.
5.1 Question Answering Performance
Among the models evaluated, the proposed Unit-LongT5-TQA exhibits remarkable performance. On the NMSQA-dev set, it achieves an F1 score of 41.8% and an EM score of 24.9%. Its performance remains consistent on the NMSQA-test dataset, registering scores of 36.0% and 24.0% for F1 and EM, respectively. Furthermore, this model outperforms the others on the generative Spoken-NarrativeQA-test set, securing a BLEU1 score of 8.0% and a ROUGE-L score of 11.8%.
5.2 The effectiveness of TQA pretraining
From Table1, we can observe that Unit-LongT5-TQA has huge improvements on both NMSQA-dev and NMSQA-test, which gets 16 F1-scores gain than Unit-LongT5, and gets two times than it on EM scores. It shows that using pretrained LM on TQA tasks as the initialization brings significant advantages for LM to learn speech information. On the other hand, in Spoken-NarrativeQA-test, an abstractive spoken QA dataset, Unit-LongT5-TQA also outperforms Unit-LongT5 on both BLEU and ROUGE scores, which gets 1.2 and 1.4 increase, respectively. In our showcase, we have observed that TQA pretraining can help the model generate more fluent sentences. The following results are from our test set. Without TQA pretraining, our prediction is “live the life of any,” where the sentence is truncated in the middle, resulting in an incoherent meaning. However, after TQA training, the model can output fluent sentences.

### 25 Dec 2023 v2

@@ Abstract
In recent advancements in spoken question answering (SQA), end-to-end models have made significant strides. However, previous research has primarily focused on extractive span selection. While this extractive-based approach is effective when answers are present directly within the input, it falls short in addressing abstractive questions, where answers are not directly extracted but inferred from the given information. To bridge this gap, we introduce the first end-to-end Generative Spoken Question Answering (GSQA) model that empowers the system to engage in abstractive reasoning. The challenge in training our GSQA model lies in the absence of a spoken abstractive QA dataset. We propose using text models for initialization and leveraging the extractive QA dataset to transfer knowledge from the text generative model to the spoken generative model. Experimental results indicate that our model surpasses the previous extractive model by 3% on extractive QA datasets. Furthermore, the GSQA model has only been fine-tuned on the spoken extractive QA dataset. Despite not having seen any spoken abstractive QA data, it can still closely match the performance of the cascade model. In conclusion, our GSQA model shows the potential to generalize to a broad spectrum of questions, thus further expanding the SQA capabilities of abstractive QA. Our code is available at https://voidful.github.io/GSQA
Index Terms—  Spoken Question Answering, Textless NLP, Generative Models, Transfer learning, Representation Learning
@@ 1 Introduction
An intuitive approach to tackle this challenge is to cascade models. Here, an Automatic Speech Recognition (ASR) system first transcribes the input speech of both questions and passages and then feeds transcriptions into a textual question answering language model that predicts the corresponding answers. Last, answers are transformed back into speech through a Text-To-Speech (TTS) model. The cascaded method’s primary advantage is its ability to harness the great power of text-based language models trained on vast textual datasets, ensuring it can tackle both extractive and abstractive QA tasks. However, the cascaded model has its drawbacks. The primary issue stems from error propagations, since the text-based language models are not typically trained on ASR-erroneous data, the presence of ASR errors can notably undermine the performance of cascaded models. Moreover, some natural languages do not possess writing text, which leads to the absence of language models that are capable of acquiring these languages. Thus, to tackle these issues, an End-to-End speech model is needed.
@@
Recently, end-to-end, textless approaches (e.g., DUAL[5]) have emerged as a solution to the cascade method’s limitations. These works [5, 6, 7] rely on the encoded representations of acoustic models, like HuBERT[8], and apply K-means clustering to transform speech representations into discrete units. These units, encapsulating both questions and passages, are fed into the model that predicts the span of the answer in the passage using start and end positions. By adopting this textless approach, DUAL effectively sidesteps the error propagation from ASR systems [9, 10, 11]. It improves the model robustness in noisy speech. However, there’s a notable limitation to the DUAL model: it is not able to handle abstractive QA scenarios, as it is designed to only output start and end positions within the input spoken passage.
As Fig.1 shows the difference between our proposed method, cascaded method, and DUAL, we combine the advantages of both the cascaded method and end-to-end (E2E) textless approach to introduce the first-ever textless E2E model capable of handling both extractive and abstractive QA. Mirroring DUAL’s methodology, we utilize HuBERT to convert the input speech (questions and passages) into discrete units[6]. A sequence-to-sequence model then processes these units to generate answers in the form of discrete units. Motivated by the strengths and limitations of the approaches above, we propose GSQA (Generative Spoken Question Answering), the first textless end-to-end model capable of handling both extractive and abstractive QA.
Furthermore, given the current absence of a spoken abstractive QA corpus, we begin by initializing GSQA with a generative Text QA (TQA) model. This TQA model has been trained on several extractive and abstractive datasets using a generative approach. Subsequently, we fine-tune the TQA model on NMSQA, a dataset dedicated to extractive spoken QA. The primary objective of GSQA is to impart the Natural Language Understanding prowess of the TQA model to GSQA. We also use Microsoft’s Azure TTS service to create a synthesized test set for evaluation. It’s noteworthy that, during the transfer learning phase, GSQA exhibited a significant enhancement, registering an 18% relative gain in the BLEU1 score in a zero-shot abstractive spoken QA setting.
@@
    Demonstrating that our proposed model has competitive performance against textless extractive models.
@@ 2 Related Work
The existing SQA methods primarily utilize Automatic Speech Recognition (ASR) [12, 13] transcripts, processed by a large language model (LLM) for QA tasks. A key challenge is that ASR errors significantly impact the LLM’s performance. For instance, less robust ASR systems produce incorrect transcripts, leading to erroneous LLM predictions. In contrast, more advanced ASR systems, such as Whisper-large with 1550M parameters, offer lower error rates and provide more accurate input for the LLM, improving QA task accuracy. However, these high-parameter models also result in longer inference times. In extractive SQA, models identify answer spans within spoken documents. DUAL [5] is noteworthy for its ability to determine these spans, but it fails when answers are not present in the input audio. TWIST [14] addresses SpeechLM’s [15] limitations, particularly its lack of semantic understanding, which often results in irrelevant or incorrect responses. While TWIST improves SpeechLM’s initialization, it still shows limited progress in transferring semantic knowledge from text to speech.
@@ 3.1 Overview & Formulation
In the realm of Question Answering (QA), we divide the problem into three components: the Question (Q), the Passage (P), and the Answer (A). These entities can manifest in two formats: textual data ($Q_{t}$, $P_{t}$, $A_{t}$) and speech discrete unit-based data ($Q_{u}$, $P_{u}$, $A_{u}$).
Our method lies in the ambition to create a textless generative end-to-end language model that is capable of question answering. Initially, we follow [6, 7] to quantize speech representations into discrete units. We then utilize a discrete unit-based sequence generative model to craft the answer units, which would be subsequently transformed into speech via a Unit-based HiFi-GAN vocoder [16]. To boost the logical understanding capabilities of the discrete unit-based sequence generative model, we initialize this generative SQA model using weights from a textual model [14]. Fig.2 shows the entire model architecture and training pipeline.
Fig. 2: Left: The process of discrete unit quantization from synthesis data. Right: Model Training Procedure: A depiction of the transition from textual QA pretraining to spoken QA fine-tuning.
@@ 3.2 Speech Quantization
HuBERT [8] is used for efficient speech representation, converting speech signals into discrete units through mask segment prediction training. This involves clustering each 20ms frame into one of K categories using k-means clustering, where the centroid ID of each cluster represents the frame’s discrete unit. We utilize the HuBERT-base model with a 100-cluster configuration, following Lee et al.’s methods [6], to achieve this conversion. Additionally, we employ run-length encoding to condense consecutive identical units into a single unit, thus reducing the length of input unit sequences.
3.3 Text Question Answering Pretraining
Due to the scarcity of training data in abstractive spoken QA, our model’s performance could be limited. To address this, we propose using a pretrained textual model to enhance semantic understanding. This model is initially trained as a generative language model and then adapted for SQA, facilitating the transfer of textual QA capabilities to speech units.
Handling the longer sequences in speech, even after deduplication, presents a challenge, as many generative textual models struggle with such long inputs [17]. Therefore, our chosen initialized model must be generative and capable of processing sequences longer than 4096 tokens [18, 19].
We select LongT5 [19] for this purpose. LongT5, an extension of the T5 encoder, incorporates global-local attention mechanisms to efficiently manage longer inputs. It combines the attention efficiency of ETC with PEGASUS’s summarization-focused pre-training, offering improved performance. We further pretrain LongT5 on textual QA datasets, which would be denoted as LongT5-TQA.
The primary goal during this phase is to train the model to process a textual question $Q_{t}$ and passage $P_{t}$ and generate the corresponding answer $A_{t}$. This pretraining is crucial, equipping our model with foundational QA skills for the subsequent phases of our training methodology.
@@ 3.4 GSQA: Generative Spoken QA Model
Fig.2 illustrates our GSQA training procedure. As mentioned in Sec.3.3, we pretrain the LongT5 on Text Question Answering (TQA) tasks to produce LongT5-TQA, which is capable of text-level extractive and abstractive QA. This model is then fine-tuned on the Extractive Spoken QA dataset, NMSQA [5]. To unify the dataset format in abstractive QA form, we modify the labels in NMSQA from time spans to the answers’ waveform. Here, the given spoken passages, questions, and answers would go through a quantization model, and would be converted into discrete units. Since LongT5-TQA does not recognize discrete units, we introduced them as new tokens. The embeddings of units are initialized by random sampling from other text token embeddings, ensuring proper tokenization and a seamless transition from text to speech. By these modifications, LongT5-TQA would be fine-tuned to predict answers in discrete units instead of text, where we denote the fine-tuned model as Unit-LongT5-TQA. Note that if the model’s denotation is not appended with TQA, it means that we don’t pretrain the model on TQA tasks. Last, to achieve speech-to-speech, we employ a HiFi-GAN vocoder [20] to convert discrete units back to speech. This vocoder includes a duration prediction module for deduplicated unit synthesis.
NMSQA-dev NMSQA-test Spoken-NarrativeQA-test Model Num. Param. F1 Score EM Score F1 Score EM Score BLEU1 ROUGE-L Cascade model (w/ ASR transcriptions) 1025M 49.1 32.0 47.3 30.4 13.5 19.9 DUAL 452M 39.4 21.9 33.6 21.2 - - Unit-LongT5 312M 25.5 12.6 20.1 9.4 6.8 10.4 Unit-LongT5-TQA (Proposed) 312M 41.8 24.9 36.0 24.0 8.0 11.8
Table 1: The table presents the speech-to-speech evaluation of our models and the number of parameters for each method. For extractive QA datasets, we utilized F1 and exact match (EM) as evaluation metrics. On the other hand, for abstractive QA datasets, we employed BLEU1[21] and Rouge-L[22] for assessment. The cascaded model in this table includes an ASR system and an LM, which is Whisper-medium.en [13] and LongT5-TQA [19], respectively. DUAL[5] is combined a HuBERT-base [8], K-means model for 128 clusters, and a Longformer [23].
@@ 4.1 Dataset
For TQA pretraining data, We utilized four extractive QA datasets: SQuAD [24], MultiRC [25], NewsQA [26], and Drop [27], along with one abstractive dataset, NarrativeQA [28], to pretrain LongT5-TQA, and use it as foundation model of GSQA. On the other hand, for SQA training, we use NMSQA [5] as our downstream dataset for training Unit-LongT5. The NMSQA is derived from SQuAD v1.1 [29] and consists of Wikipedia paragraphs with human-written questions. While its train and dev sets feature synthesized speech using Amazon Polly, the test set contains audio from 60 human speakers (30 males and 30 females). In addition, we propose a new spoken abstractive SQA evaluation set, named Spoken-NarrativeQA. This dataset is built upon the NarrativeQA [28] test set, an abstractive text question answering dataset for reading comprehension from books and movie scripts with complex narrative understanding. Note that many answers in this dataset are not directly extractable from passages. We transform the text to speech using the TTS system and specifically select test data where answers are absent in the passage. There are 1626 testing samples in the Spoken-NarrativeQA dataset.
@@
Table 2: The respective number of parameters in our experiment. Note that the E2E method, like DUAL and our work, does not need an ASR system but needs a Unit-Extraction model to generate units, and the number after the name of that is the number of clusters.
@@
4.2 Cascaded Pipeline
The conventional approach for SQA tasks typically follows a cascaded pipeline, as illustrated in the middle of Fig. 1. The passages and questions are first transcribed by an ASR system, followed by a text-based language model that generates text-based answers, and a TTS system converts the text-based answers into speech answers. To be more specific, we use Whisper-medium.en [13] as the ASR system, the fine-tuned LongT5-TQA model as the text-based language model, and the Azure TTS service as the TTS system. We record the experiment result of the cascaded pipeline in the column of Cascade Model in Table 1. The total number of parameters of the cascaded model is shown in Table 2, and the WERs of the ASR system in given datasets are listed in Table 3.
4.3 Automatic Performance Evaluation
Since all our models ultimately output in speech, what we focus on is whether the content of this waveform contains the answer. We then use an ASR model to convert the generated waveform back to text, allowing us to use text-based metrics to evaluate the content of the model. To reduce errors from the ASR during evaluation, we utilize the state-of-the-art ASR model, whisper-large-v2, during automatic performance evaluation. We then compare the transcriptions of the outputs from whisper-large-v2 with the text ground truth. We calculate the F1-score and Exact-matched (EM) scores for extractive QA tasks. The BLEU1 score [21] and the Rouge score [22] are used for abstractive QA tasks. Lastly, the experiment results of all models are reported in Table 1.
4.4 Implementation details
As the previous section mentioned, we pretrain LongT5 on the 5 TQA datasets for 13 epochs, the pretraining learning rate is set to 0.0005, and the weight decay is 0.01. On the other hand, we fine-tune the pretrained LongT5 on the downstream unit-based dataset, NMSQA, for 25 epochs, the downstream learning rate is set to 0.0003, and the weight decay is 0.001. Last, when we inference our model on the abstractive test set(e.g., Spoken-NarrativeQA-test), to encourage GSQA output longer answers, we use beam search for decoding with beam size of 5, and set length penalty to 2.
5 Results
As illustrated in Table 1, we evaluate our models on both extractive QA dataset NMSQA-dev, NMSQA-test, and on abstractive QA dataset Spoken-Narrative QA. Among the models evaluated, the proposed Unit-LongT5-TQA exhibits remarkable performance. On the NMSQA-dev set, it achieves an F1 score of 41.8% and an EM score of 24.9%. Its performance remains consistent on the NMSQA-test dataset, registering scores of 36.0% and 24.0% for F1 and EM, respectively. Furthermore, this model outperforms the others on the generative Spoken-NarrativeQA-test set, securing a BLEU1 score of 8.0% and a ROUGE-L score of 11.8%.
5.1 The effectiveness of TQA pretraining
From Table1, we can observe that Unit-LongT5-TQA has huge improvements on both NMSQA-dev and NMSQA-test, which gets 16% F1-scores gain than Unit-LongT5 and gets two times than it on EM scores. It shows that pretraining language model on TQA tasks brings significant advantages for unit-based model initialization to learn semantic information in speech. On the other hand, in the Spoken-NarrativeQA-test, an abstractive spoken QA dataset, Unit-LongT5-TQA also outperforms Unit-LongT5 on both BLEU1 and ROUGE-L scores, which gets 1.2 and 1.4 increase, respectively. In our showcase, we have observed that TQA pretraining could further help the model generating more fluent sentences. The following results in Table 4 are from our test set. Without TQA pretraining, our prediction is ”live the life of any,” where the sentence is truncated in the middle, resulting in an incoherent meaning. However, after TQA training, the model would predict complete sentences.